OpenStudy (anonymous):

PLEASE IF U CAN HELP!!!! I NEED HELP FAST!!!! Performance Task: You have been asked to prepare a demonstration in which you will analyze the act of dropping the watermelon from a height of 160 feet with an initial speed of 48 ft/sec. The height is modeled by s(t) = -16t2 – 48t + 160. Your report needs to include answers to the following: ; When will the watermelon strike the ground? • What is the average speed of the watermelon from t = 0 to t = 1? What is the instantaneous speed of the watermelon at time t = 1? What is the instantaneous speed of the watermelon when it strikes the ground?

OpenStudy (agent47):

the watermelon will strike the ground when: the height, or s(t)=0.

OpenStudy (agent47):

-16t2 - 48t + 160 = 0 t=-5 t=2 Time can't be negative so the watermelon strikes the ground at t = 2 seconds

OpenStudy (agent47):

Average speed is distance traveled over the time it took to travel: [s(1)-s(0)]/(1-0)=(96-160)/1=-64

OpenStudy (agent47):

-64ft/sec that is. Sorry i'm kinga slow today, haven't done any math in a while.

OpenStudy (anonymous):

ur just fine thank you for helping me :)

OpenStudy (agent47):

What is the instantaneous speed of the watermelon at time t = 1? V(t)=-32t-48 V(1)=-32*1-48=-80 ft/sec

OpenStudy (agent47):

What is the instantaneous speed of the watermelon when it strikes the ground? It strikes the ground at t=2 V(2)=-32*2-48=-64-48=-112